5) Michigan Company has budgeted the following costs for the production of its only product:

Direct Materials                                                     $35,000

Direct Labor                                                           25,000

Variable indirect production costs                            30,000

Fixed indirect production costs                                15,000

Variable selling and administrative costs                    7,500

Fixed selling and administrative costs                      12,500

Total Costs                                                          $125,000

Michigan Company wants a profit of $50,000, and expects to produce 1,000 units. The market price is $150 per unit. What is the target cost per unit of the product?

A) $100 per unit (this is the right answer) PLEASE EXPLAIN WHY

B) $125 per unit

C) $150 per unit

D) $175 per unit
Answer #1

Profit needed per unit = 50,000/1,000 = 50 per unit

Target selling price per unit = Target cost per unit + Target profit per unit

150 = Target cost per unit + 50

Target cost per unit = 150 - 50

= $100